Suppose that a certain market supplies x thousands of crates of orange daily when p dollars is the price per crate, and the supply equation is px-20p-3x+105=0. If the daily supply decreases at the rate of 250 crates per day, at what rate is the price changing when the daily supply is 5000 crates? 


1
Expert's answer
2021-07-26T16:33:15-0400
"px-20p-3x+105=0"

Differentiate with respect to "t"


"p'x+px'-20p'-3x'=0"

Solve for "p'"


"p'=\dfrac{(3-p)x'}{x-20}"

Given "x'=-0.250, x=5"


"p(5)-20p-3(5)+105=0"

"p=6"

Substitute


"p'=\dfrac{(3-6)(-0.25)}{5-20}=-0.05"

The price is dropping by a nickel ($0.05).